Find the p -value for the hypothesis test. A random sample of size 52 is taken. The sample has a mean of 388 and a standard deviation of 85. H0: µ = 400 Ha: µ< 400 The p -value for the hypothesis test is?
Answer #1

Solution:

Given:

Sample size = n= 52

Sample mean = \bar{x} = 388

Sample standard deviation = s = 85

H0: µ = 400 Ha: µ< 400

We have to find the p -value for the hypothesis test.

Since sample size =n =52 is large ,we will use z test statistic and sample standard deviation as an estimate of population standard deviation

thus find z test statistic:
z=\frac{\bar{x}-\mu }{\sigma /\sqrt{n}}

z=\frac{ 388 - 400 }{85 /\sqrt{52}}

z=\frac{ -12 }{85 /7.211103 }

z=\frac{ -12 }{11.7873792 }

z=-1.02

Thus p-value is:

p-value = P(Z < z test statistic)

p-value = P(Z < -1.02)

Look in z table for z = -1.0 and 0.02 and find corresponding area.

.00 01 .02 .03 .04 .05 .06 07 .08 .09 z .0003 .0005 .0006 .0009 .0013 .0018 .0024 .0033 .0044 .0059 .0078 .0102 .0 32 0003 00

Thus

P( Z < -1.02 ) = 0.1539

thus

p-value = P(Z < -1.02)

p-value = 0.1539